# S2CFormer: Reorienting Learned Image Compression from Spatial Interaction to Channel Aggregation

### v2

@@ Abstract
Transformers have achieved significant success in learned image compression (LIC), with Swin Transformers emerging as the mainstream choice for nonlinear transforms. A common belief is that their sophisticated spatial operations contribute most to their efficacy. However, the crucial role of the feed-forward network (FFN) based Channel Aggregation module within the transformer architecture has been largely overlooked, and the over-design of spatial operations leads to a suboptimal trade-off between decoding latency and R-D performance. In this paper, we reevaluate the key factors behind the competence of transformers in LIC. By replacing spatial operations with identity mapping, we are surprised to find that channel operations alone can approach the rate-distortion (R-D) performance of the leading methods. This solid lower bound of performance emphasizes that the presence of channel aggregation is more essential for the LIC model to achieve competitive performance, while the previously complex spatial interactions are partly redundant. This motivates us to use simplified spatial operations and powerful channel operations to overcome the previous trade-off. Based on this insight, we initiate the “S2CFormer” paradigm, a general architecture that reorients the focus of LIC from Spatial Interaction to Channel Aggregation. We present two instantiations of the S2CFormer: S2C-Conv, and S2C-Attention. Each one incorporates a simple operator for spatial interaction and serves as nonlinear transform blocks for our LIC models. Both models demonstrate state-of-the-art (SOTA) R-D performance and significantly faster decoding speed, as shown in Fig. 1. These results also motivate further exploration of advanced FFN structures to enhance the R-D performance while maintaining model efficiency. With these foundations, we introduce S2C-Hybrid, an enhanced LIC model that combines the strengths of different S2CFormer instantiations. This model outperforms all the existing methods on the Kodak, Tecnick, and CLIC Professional Validation datasets, setting a new benchmark for efficient and high-performance LIC. The code is at https://github.com/YunuoChen/S2CFormer.
Figure 1: S2CFormer and the performance of S2CFormer-based models. The general structure of our S2CFormer is shown in (a). It consists of two key components: the Spatial Interaction module and the Channel Aggregation module. S2CFormer functions as nonlinear transform blocks for Learned Image Compression (LIC). We argue that the competence of transformer-based LIC models primarily stems from the Channel Aggregation module, rather than the specific Spatial Interaction module. As illustrated in (b), S2CFormer-based models not only outperform leading LIC models in rate-distortion performance but also offer significant improvements in decoding latency.
@@ 1 Introduction
The success of transformers, particularly Swin Transformers, is often attributed to their sophisticated spatial operations, such as self-attention and window shifting. However, the significance of the transformer structure has not been thoroughly explored, making such attributions premature. Besides, their heavy reliance on spatial operations leads to a suboptimal trade-off between decoding latency and R-D performance.
@@
To this end, we reevaluate the key factors in the R-D performance of transformers-based LIC models. We identify that a basic vision transformer block [47, 33, 34] consists of two operations targeting different dimensions: an attention-based spatial operation and an MLP-based channel operation. By replacing spatial operations with identity mapping, we are surprised to find that channel operations alone can achieve promising performance, which is comparable to the leading methods. The solid lower performance bound and the simple identity spatial operation emphasize that the MLP-based channel operation plays an essential role in optimizing R-D performance for LIC models, while previously complex spatial interactions are partly redundant. This result also gives us hope of overcoming the trade-off between R-D performance and decoding latency inherent in previous methods. To reorient the focus of LIC from Spatial Interaction to Channel Aggregation, we introduce the paradigm “S2CFormer”. S2CFormer is a general structure originated from the standard vision transformers [47, 33, 34], which consists of two main components: Spatial Interaction and Channel Aggregation. This paradigm emphasizes two key ideas: 1) simplifying spatial interactions to accelerate decoding, and 2) aggregating anisotropic features along the channel dimension to achieve promising performance. It realigns the priority of the spatial operator and channel operator to break the limitations of the previous speed-performance trade-off.
@@
Based on the insight of S2CFormer, we adopt two basic operators for spatial interaction: separable convolution, and vanilla attention. These operators, combined with an MLP-based FFN for channel aggregation, constitute two basic S2CFormer instantiations: S2C-Conv, and S2C-Attention. S2C-Conv and S2C-Attention demonstrate that combining simple spatial interaction with channel aggregation can achieve SOTA performance and improve decoding speed by over 30%. Their performances indicate that channel aggregation is the key factor in R-D performance, far outweighing the choice of specific spatial interaction methods (such as convolution or attention). We suppose this is due to the fact that as the resolution of the feature decreases, spatial redundancy shifts to channels, making it crucial to capture and eliminate channel redundancy.
@@
These findings motivate further exploration of advanced FFN structures for channel aggregation, which could enhance LIC models with minimal added complexity. Additionally, utilizing the advantages of different S2C-instantiations, we propose a new LIC model, S2C-Hybrid. This model combines the strengths of both attention and convolution mechanisms for spatial interaction, while employing the advanced Gated-FFN for channel aggregation, thereby significantly improving R-D performance. Besides, it also ensures a much faster decoding speed compared to existing SOTA methods.
@@
    Establishing New Performance Benchmarks: The S2C-Hybrid model combines the strengths of different S2C-instantiations and outperforms all the existing methods on three datasets (i.e., Kodak, Tecnick, and CLIC datasets), establishing a new R-D performance record for LIC.
@@ 2.1 Learned Image Compression
In the past few years, end-to-end learned image compression (LIC) has gained increased attention. Ballé et al. [2] proposed the first end-to-end image compression model with convolution neural network (CNN) and introduced the variation auto-encoder (VAE) model combined with hyper-prior [3], which has become the most fundamental structure for LIC models. Afterward, researchers mainly focus on two things to improve the rate-distortion performance: transform networks and entropy model [52, 39, 37, 35, 31, 17, 16, 4].
@@ 2.2 Transformer
Transformers, initially introduced by [47] for translation tasks, has quickly become widely adopted in a range of NLP applications. Building on this success, many researchers have expanded the use of attention mechanisms and transformers to tackle computer vision tasks. [15, 33, 46, 53, 9].
In particular, Liu et al. [33] introduced the Swin Transformer, a hierarchical model that employs window attention and window shifting mechanism. This innovative structure has established a new performance benchmark in both high-level [48, 14, 34] and low-level [30, 53, 9, 54] vision tasks, reflecting its broad applicability and effectiveness.
Many learned image compression models have adopted the Swin Transformer for nonlinear transform, achieving remarkable improvements in rate-distortion performance. The success of those advanced LIC models has been long attributed to the sophisticated attention modules. However, in the field of natural language processing (NLP), Thorp et al. suggest in their work [26] that substituting the attention module with a simple Fourier transform can yield comparable results. Similarly, in high-level vision tasks, Yu et al. [50, 51] demonstrate that the architecture known as MetaFormer is actually the most crucial factor for achieving high performance.
In this paper, we revisit the competence of transformers in LIC models, arguing that the core factor for superior R-D performance is Channel Aggregation, rather than the well-designed attention modules.
@@ 3.1 Problem Formulation
Figure 2: Overview of S2CFormer-based LIC model. We adopt the basic VAE structure from [41, 3] and integrate the SCCTX entropy model from [20]. The hierarchical architecture consists of five stages of nonlinear transform blocks. Each stage contains $L_{i}$ S2CFormer blocks. The general S2CFormer architecture is shown in (a), and (b-d) illustrate three S2CFormer instantiations. $L_{1}$-$L_{6}$ represent block numbers for each stage.
The structure of our S2CFormer-based LIC model is illustrated in Fig. 2. It consists of three main components: an Encoder, a Decoder, and an Entropy Model. The Encoder generates a latent representation $\bm{y}$ from RGB input $\bm{x}$, which is quantized to $\hat{\bm{y}}$. The Decoder maps $\hat{\bm{y}}$ back to the RGB domain. This process involves an encoder network $g_{a}(\cdot)$, a decoder network $g_{s}(\cdot)$, and a quantization operator $Q$. The steps are:
@@
We focus on the design of four transform networks: Encoder $g_{a}$, Decoder $g_{s}$, Hyper Encoder $h_{a}$, and Hyper Decoder $h_{s}$, each including several nonlinear transform blocks. The S2CFormer block is employed for these transforms. Downsampling is achieved with a $3\times 3$ convolution layer (stride 2), and upsampling uses a pixel shuffle layer with a $3\times 3$ convolution. The entropy models adopt the efficient Space-Channel ConTeXt (SCCTX) model from ELIC [20], enhancing entropy estimation with spatial and channel context.
@@ 3.2.1 S2C-Identity: The Lower Bound of S2CFormers
Firstly, we introduce the “S2CFormer” paradigm, shown in Fig. 2 (a). S2CFormer is a general architecture comprising two main components: Spatial Interaction and Channel Aggregation.
@@ 3.2.2 Simplified Spatial Interaction
In addition to S2C-Identity, we explore more S2CFormer instantiations with other spatial operators, as shown in Fig. 2 (c-d). Instead of designing novel spatial interaction modules, we assess the model’s potential for achieving SOTA performance with simplified, commonly used spatial operators. We implement two spatial interaction modules. The first is separable convolution, as proposed in [50, 11, 45]:
@@ 3.2.3 Advanced Channel Aggregation
These advanced FFNs outperform the vanilla FFN while maintaining similar computational complexity and latency. This suggests that exploring efficient FFN designs is crucial for effective LIC models, beyond relying solely on complex spatial interactions.
3.2.4 S2C-Hybrid: A Superior Extension of S2CFormers
Building on the spatial interaction operators described in Sec. 3.2.2 and the advanced channel aggregation techniques from Sec. 3.2.3, we introduce the S2C-Hybrid model, an enhanced version of the S2CFormer-based LIC model. S2C-Hybrid offers an excellent trade-off between R-D performance and decoding speed.
This hybrid model employs different S2C instantiations for different transform stages. For channel aggregation, we adopt gated-FFN to enhance performance. As for spatial interaction, S2C-Hybrid takes advantage of both convolutional networks and self-attention mechanisms. Specifically, we employ convolution in the early stages when feature resolution is high, and apply attention mechanisms in the later stages. This strategy improves R-D performance without significantly increasing time complexity or slowing down coding speed. Additionally, the S2C-Conv and S2C-Identity blocks are used to enhance spatial context extraction and parameter aggregation in the SCCTX entropy model.
With advanced channel aggregation to enhance performance and simplified spatial interactions to accelerate speed, S2C-Hybrid overcomes the suboptimal trade-off between performance and speed inherent in previous methods.
ELIC MambaVC FTIC MLIC++ TCM-large S2C-Identity S2C-Conv S2C-Attention S2C-Hybrid
BD-rate [Kodak] -3.10 -8.11 -12.94 -11.97 -10.04 -9.70 -12.65 -12.56 -14.28
BD-rate [CLIC] -0.84 - -10.21 -12.08 -8.60 -7.55 -10.98 -10.28 -12.88
BD-rate [Tecnick] -7.41 - -13.89 -15.13 -10.42 -10.70 -14.48 -14.08 -17.20
Params (M) 33.29 47.88 69.78 116.48 75.89 64.63 66.60 68.42 79.83
FLOPs (T) 1.74 2.10 2.38 2.64 3.74 2.80 3.13 3.42 4.17
@@ 3.2.4 S2C-Hybrid: A Superior Extension of S2CFormers
Throughput (samples/s) $\uparrow$ 72.32 6.55 23.25 27.42 17.80 50.02 35.49 26.21 22.63
Table 1: Comprehensive comparisons of LIC models. (a) BD-rate. BD-rate of LIC models on different datasets, relative to VTM-21.0 [7]. S2C-Conv, -Attention, and -Hybrid models achieve superior R-D performance. R-D curves are provided in Fig. 6-8. (b) Model Params. S2C-Conv and -Attention models have fewer parameters compared to previous SOTA methods. (c) Efficiency. FLOPs assess computational complexity, but they do not always reflect the actual training and inference speed. While S2CFormer-based models do not exhibit many advantages in terms of FLOPs, they achieve over 30% faster decoding speed and significantly higher training throughput.
@@ 4.1.1 Training Details
We train all S2CFormer-based models on the Flickr2W dataset [31] for 2 million steps with a batch size of 8, using Adam [23] optimizer and an initial learning rate of 0.0001. For MSE-optimized models, we use Lagrangian multipliers {0.0017, 0.0025, 0.0035, 0.0067, 0.0130, 0.0250, 0.050}; for MS-SSIM-optimized models, we use {3, 5, 8, 16, 36, 64}. In the attention modules, the window size is set to 8, and depthwise convolution kernels are all of size 5. The latent representation $y$ consists of 320 channels, and the hyper-latent representation contains 192 channels. Additional details are provided in the supplementary material. To enhance context mining and parameter aggregation within the entropy model of ELIC [20], we incorporate the S2C-Conv block and the S2C-Identity block, respectively. Unless otherwise specified, we use the Gated-FFN as a channel aggregation module for all S2CFormer-based models. All experiments are conducted on NVIDIA A100 GPUs.
@@ 4.1.2 Evaluation
4.2 Rate-Distortion Performance
In this section, we compare our models with VTM-21.0 [7, 6] and several leading LIC models, including ELIC [20], MambaVC [43], TCM [32], MLIC++ [22] and FTIC [27]. The detailed results are presented in Table 1.

### v3

# S2CFormer: Revisiting the RD-Latency Trade-off in Transformer-based Learned Image Compression

@@ Abstract
Transformer-based Learned Image Compression (LIC) suffers from a suboptimal trade-off between decoding latency and rate-distortion (R-D) performance. Moreover, the critical role of the FeedForward Network (FFN)-based channel aggregation module has been largely overlooked. Our research reveals that efficient channel aggregation—rather than complex and time-consuming spatial operations—is the key to achieving competitive LIC models. Based on this insight, we initiate the “S2CFormer” paradigm, a general architecture that simplifies spatial operations and enhances channel operations to overcome the previous trade-off. We present two instances of the S2CFormer: S2C-Conv, and S2C-Attention. Both models demonstrate state-of-the-art (SOTA) R-D performance and significantly faster decoding speed, as shown in Fig. 1. Furthermore, we introduce S2C-Hybrid, an enhanced variant that maximizes the strengths of different S2CFormer instances to achieve a better performance-latency trade-off. This model outperforms all the existing methods on the Kodak, Tecnick, and CLIC Professional Validation datasets, setting a new benchmark for efficient and high-performance LIC. The code is at https://github.com/YunuoChen/S2CFormer.
Figure 1: S2CFormer and the performance of S2CFormer-based models. The general structure of our S2CFormer is shown in (a). It consists of two key components: the Spatial Interaction module and the Channel Aggregation module. S2CFormer functions as nonlinear transform blocks for Learned Image Compression (LIC). Our analysis reveals that the competence of transformer-based LIC models primarily stems from channel aggregation. Building on this insight, we propose a novel design strategy that rebalances these two modules to achieve a more favorable trade-off between compression performance and decoding latency. As illustrated in (b), the data points for S2CFormer-based models exhibit a linear trend with a steeper slope, thereby underscoring their superior performance–latency characteristics.
@@ 1 Introduction
Figure 2: Comparison of execution times for spatial interaction and channel aggregation across different models. Previous methods show much higher spatial interaction times than channel aggregation, causing significant delays. Our S2CFormer effectively rebalances the time relationship between these two modules.
@@
The success of transformers is often attributed to their sophisticated spatial operations, such as self-attention and window shifting. However, the significance of the transformer structure has not been thoroughly explored, making such attributions premature. Besides, as shown in Fig. 1 and Fig. 2, their complex and heavy spatial operations are quite time-consuming, thus often leading to a suboptimal trade-off between decoding latency and R-D performance.
@@
To this end, we reevaluate the key factors in the R-D performance of transformers-based LIC models. We identify that a basic vision transformer block [48, 34, 35] consists of two operations targeting different dimensions: an attention-based spatial operation and an MLP-based channel operation. By replacing spatial operations with identity mapping, we are surprised to find that channel operations alone can achieve promising performance, which is even comparable to the leading methods. The solid lower performance bound and the simple identity spatial operation emphasize that the MLP-based channel operation plays an essential role in optimizing R-D performance for LIC models, while previously complex spatial interactions may be partly redundant. To reorient the focus of LIC from Spatial Interaction to Channel Aggregation, we introduce the paradigm “S2CFormer”. S2CFormer is a general structure originated from the standard vision transformers, which consists of two main components: Spatial Interaction and Channel Aggregation. This paradigm emphasizes two key ideas: 1) simplifying spatial interactions to accelerate decoding, and 2) aggregating anisotropic features along the channel dimension to achieve promising performance. It realigns the priority of the spatial operator and channel operator to break the limitations of the previous speed-performance trade-off.
@@
Based on the insight of S2CFormer, we adopt two basic operators for spatial interaction: separable convolution, and vanilla attention. These operators, combined with a FFN for channel aggregation, constitute two basic S2CFormer instances: S2C-Conv, and S2C-Attention. S2C-Conv and S2C-Attention demonstrate that combining simple spatial interaction with channel aggregation can achieve SOTA performance and improve decoding speed by over 30%. Their performances indicate that channel aggregation is the key factor in R-D performance, far outweighing the choice of specific spatial interaction methods (e.g., convolution or attention). We attribute this to reduced feature resolution shifting spatial redundancy to channels, highlighting the importance of capturing and eliminating channel redundancy.
@@
These findings motivate further exploration of advanced MLP structures for channel aggregation, which could enhance LIC models with minimal added complexity. Additionally, we propose a new LIC model, S2C-Hybrid. This model achieves even better compression performance by appropriately arranging and combining different S2C-instances, thus fully exploiting the strengths of each S2C-instantiation. Besides, it also ensures a much faster decoding speed compared to existing SOTA methods.
@@
    Establishing New Performance Benchmarks: The S2C-Hybrid model combines the strengths of different S2C-instances and outperforms all the existing methods on three datasets (i.e., Kodak, Tecnick, and CLIC datasets), establishing a new R-D performance record for LIC.
@@ 2.1 Learned Image Compression
In the past few years, end-to-end learned image compression (LIC) has gained increased attention. Ballé et al. [2] proposed the first end-to-end image compression model with convolution neural network (CNN) and introduced the variation auto-encoder (VAE) model combined with hyper-prior [3], which has become the fundamental structure for LIC models. Afterward, researchers mainly focus on two things to improve the rate-distortion performance: transform networks and entropy model [53, 40, 38, 36, 32, 17, 16, 4, 18].
@@ 2.2 Transformer
Transformers, initially introduced by [48] for translation tasks, has quickly become widely adopted in a range of NLP applications. Building on this success, many researchers have expanded the use of attention mechanisms and transformers to tackle computer vision tasks. [15, 34, 47, 54, 9]. In particular, Liu et al. [34] introduced the Swin Transformer, a hierarchical model that employs window attention and window shifting mechanisms. This innovative structure has established a new performance benchmark in both high-level [49, 14, 35] and low-level [31, 54, 9, 55] vision tasks.
Many LIC models have adopted the Swin Transformer for nonlinear transform, achieving remarkable improvements in RD performance. The success of those models has been long attributed to the sophisticated attention modules. However, in the field of natural language processing (NLP), Thorp et al. suggest in their work [27] that substituting the attention module with a simple Fourier transform can yield comparable results. Similarly, in high-level vision tasks, Yu et al. [51, 52] demonstrate that the architecture known as MetaFormer is actually the most crucial factor for achieving high performance. In this paper, we revisit the competence of transformers in LIC models, arguing that the core factor for superior R-D performance is Channel Aggregation, rather than the well-designed attention modules.
@@ 3.1 Problem Formulation
Figure 3: Overview of S2CFormer-based LIC model. We adopt the basic VAE structure from [42, 3] and integrate the SCCTX entropy model from [21]. The hierarchical architecture consists of five stages of nonlinear transform blocks. Each stage contains $L_{i}$ S2CFormer blocks. The general S2CFormer architecture is shown in (a), and (b-d) illustrate three S2CFormer instances. $L_{1}$-$L_{6}$ and $C_{1}$-$C_{6}$ represent block numbers and channel numbers for each stage, respectively
The structure of our S2CFormer-based LIC model is illustrated in Fig. 3. It consists of three main components: an Encoder, a Decoder, and an Entropy Model. The Encoder generates a latent representation $\bm{y}$ from RGB input $\bm{x}$, which is quantized to $\hat{\bm{y}}$. The Decoder maps $\hat{\bm{y}}$ back to the RGB domain. This process involves an encoder network $g_{a}(\cdot)$, a decoder network $g_{s}(\cdot)$, and a quantization operator $Q$:
@@ 3.2 S2CFormer
In this paper, we focus on overcoming the previously sub-optimal RD-latency trade-off by rebalancing spatial interaction and channel aggregation. Existing studies have typically attributed the success of transformers in RD performance to attention-based spatial interactions, while research on channel aggregation modules for LIC remains limited. However, as illustrated in Fig. 2, excessively complex spatial interaction modules significantly degrade decoding speed, whereas channel aggregation appears comparatively more efficient.
Consequently, we aim to determine whether channel aggregation might be the primary factor affecting RD performance. To this end, we propose a new paradigm called “S2CFormer,” which reorients LIC from spatial interaction towards channel aggregation by significantly simplifying spatial interaction modules. As depicted in Fig. 3 (a), our proposed paradigm is a general architecture consisting of two primary components: Spatial Interaction and Channel Aggregation.
@@ 3.2.2 Simplified Spatial Interaction
In addition to S2C-Identity, we explore more S2CFormer instances with other spatial operators, as shown in Fig. 3 (c-d). Instead of designing novel spatial interaction modules, we assess the model’s potential for achieving SOTA performance with simplified, commonly used spatial operators. We implement two spatial interaction modules. The first is separable convolution, as proposed in [51, 11, 46]:
@@ 3.2.3 Advanced Channel Aggregation
Advanced FFNs surpass vanilla FFNs in performance with comparable computational complexity and latency, highlighting the importance of efficient FFN designs for effective LIC models, beyond complex spatial interactions.
3.2.4 S2C-Hybrid: A Superior Extension.
The strong RD performance of S2C-Conv and S2C-Attention highlights the redundancy of prior complex spatial operations but raises new questions: While the LIC model combining these modules achieves comparable overall performance, do they perform equally at each stage? If not, could an optimal configuration maximize their strengths at specific stages, further improving LIC performance without sacrificing efficiency? Through experiments, we identified the best configuration under the current framework, with detailed results in Sec. 4.4.
On such basis, we introduce the S2C-Hybrid model, an enhanced version of the S2CFormer-based LIC model. This hybrid model employs different S2C instances for different transform stages. Specifically, we employ convolution in the first stage when feature resolution is high, and apply attention mechanisms in the later two stages. This strategy improves R-D performance without increasing time complexity or slowing down coding speed. S2C-Hybrid offers an even more excellent trade-off between R-D performance and decoding speed.
Previous SOTA Methods S2C Basic Instances S2C-Hybrid and Scaling-Up
ELIC MambaVC FTIC MLIC++ TCM S2C-Identity S2C-Conv S2C-Attention Hybrid-S Hybrid-M Hybrid-L
BD-rate [Kodak] -3.10 -8.11 -12.94 -11.97 -10.04 -9.70 -12.65 -12.56 -13.35 -13.85 -14.28
BD-rate [CLIC] -0.84 - -10.21 -12.08 -8.60 -7.55 -10.98 -10.28 -12.07 -12.52 -12.88
BD-rate [Tecnick] -7.41 - -13.89 -15.13 -10.42 -10.70 -14.48 -14.08 -15.21 -16.34 -17.20
Params (M) 33.29 47.88 69.78 116.48 75.89 64.63 66.60 68.42 68.00 72.73 79.83
FLOPs (T) 1.74 2.10 2.38 2.64 3.74 2.80 3.13 3.42 3.20 3.59 4.17
@@
Throughput(samples/s) 72.32 6.55 23.25 27.42 17.80 50.02 35.49 26.21 30.21 28.07 22.63
Table 1: Comprehensive comparisons of LIC models. (a) BD-rate. BD-rate of LIC models on different datasets, relative to VTM-21.0 [7]. S2C-Conv, -Attention, and -Hybrid models achieve superior R-D performance. (b) Model Params. S2C-based models have fewer parameters compared to previous SOTA methods. (c) Efficiency. S2C-based models achieve over 30% faster decoding speed and significantly higher training throughput. However, they are not superior in terms of FLOPs, and we will explain this issue in 4.5.
@@ 4.1.1 Training Details
We train all S2CFormer-based models on the Flickr2W dataset [32] for 2 million steps with a batch size of 8, using Adam [24] optimizer and an initial learning rate of 0.0001. For MSE-optimized models, we use Lagrangian multipliers {0.0017, 0.0025, 0.0035, 0.0067, 0.0130, 0.0250, 0.050}; for MS-SSIM-optimized models, we use {3, 5, 8, 16, 36, 64}. In the attention modules, the window size is set to 8, and depthwise convolution kernels are all of size 5. We set {192, 192, 192, 320} for channel numbers $C_{1},C_{2},C_{3},C_{4}$. To enhance context mining and parameter aggregation within the entropy model of ELIC [21], we incorporate the S2C-Conv block and the S2C-Identity block, respectively. Unless otherwise specified, we use the Gated-FFN as a channel aggregation module for all S2CFormer-based models. All experiments are conducted on NVIDIA A100 GPUs.
@@ 4.1.2 Evaluation
We compare our models with VTM-21.0 [7, 6] and several leading LIC models, including ELIC [21], MambaVC [44], TCM [33], MLIC++ [23] and FTIC [28]. The detailed results are presented in Table 1.
4.2 Basic S2CFormer Instances